OpenStudy (anonymous):

9x^2 + 25y^2 = 225 The foci are: (-4, 0) and (4, 0) (0, -4) and (0, 4) (-5, 0) and (5, 0)

OpenStudy (xapproachesinfinity):

hmm how about you start by dividing by 225

OpenStudy (xapproachesinfinity):

\[\frac{9}{225}x^2+\frac{25}{225}y^2=1\]

OpenStudy (xapproachesinfinity):

hmm willing to work?

OpenStudy (xapproachesinfinity):

i need you to work out a little bit, reduce those two fractions

OpenStudy (anonymous):

If you're going to be rude, then no I am not willing to work with YOU.

OpenStudy (triciaal):